Lynee is eight years older than Cherese. Four years ago she was twice as old as Cherese was. How old is Lynne now?

Answer by stanbon(75887)

Lynee is eight years older than Cherese.
Four years ago she was twice as old as Cherese was.
How old is Lynne now?
-----
Equations:
L = C + 8
(L-4) = 2(C-4)
--------------------
Substitute for "L" and solve for "C":
(C+8-4) = 2C-8
C+4 = 2C-8
C = 12 yrs (Cherese's age now)
---
L = C+8 = 20 yrs (Lynee's age now)
======================================
Cheers,
Stan H.
==========